# Optimal Computation in Leaderless and Multi-Leader Disconnected Anonymous Dynamic Networks

### 20 Feb 2024 v5

@@ 32   Return $(n^{\prime},t)$
Note that the assumptions of Lemma A.4 are satisfied, and therefore all the conditional anonymities and weights assigned to nodes up to this point are correct according to our definitions.
We will begin by proving that, if the subtree hanging from a node $v$ of $\mathcal{V}$ contains more than $a^{\prime}(v)$ nodes marked as guessed, then it contains a node $v^{\prime}$ marked as guessed such that $w(v^{\prime})>a^{\prime}(v^{\prime})$. The proof is by well-founded induction based on the subtree relation in $\mathcal{V}$. If $v$ is guessed, then we can take $v^{\prime}=v$. Otherwise, by the pigeonhole principle, $v$ has at least one child $u$ whose hanging subtree contains more than $a^{\prime}(u)$ guessed nodes. Thus, $v^{\prime}$ is found in this subtree by the induction hypothesis.
Assume for a contradiction that more than $\delta n$ levels of $\mathcal{V}$ are locked; hence, $\mathcal{V}$ contains more than $\delta n$ nodes marked as guessed. Since the conditional anonymity of the root $r$ of $\mathcal{V}$ is $\delta n$, by the above paragraph we know that $\mathcal{V}$ contains a guessed node $v^{\prime}$ such that $w(v^{\prime})>a^{\prime}(v^{\prime})$. Since $\delta\geq 1$ and Lemma A.4 (i) holds, we can apply Lemma A.2 to $v^{\prime}$, which implies that there exist heavy nodes. In turn, this contradicts Lemma A.4 (ii). We conclude that at most $\delta n$ levels are locked. ∎
We say that a node $v$ of the history tree $\mathcal{H}$ is missing from level $L_{i}$ of the view $\mathcal{V}$ if $v$ is at the level of $\mathcal{H}$ corresponding to $L_{i}$ but does not appear in $\mathcal{V}$. Clearly, if a level of $\mathcal{V}$ has no missing nodes, all previous levels also have no missing nodes.
Lemma A.6.
Assume that $\delta\geq 1$. Then, as long as level $L_{t}$ of $\mathcal{V}$ is not missing any nodes (where $t$ is defined and updated as in ApproxCount), whenever Line 13 is reached, there are at most $n-2$ levels in the range from $L_{s+1}$ to $L_{t}$ that lack a guessable non-counted node.
@@
By definition of $t$, either $t=s$ or the algorithm has performed at least one guess on a node at level $L_{t}$ with a guesser at level $L_{t-1}$. It is easy to prove by induction that the first guesser to perform a guess on this level must be the unique descendant $\tau^{\prime}\in L_{t-1}$ of the selected leader node $\tau\in L_{s}$. Moreover, both $\tau^{\prime}$ and its unique child in $\mathcal{V}$ have been assigned conditional anonymity $x$ at Lines 9–11, and the same is true of all nodes in the black path $P_{\tau}$ from $\tau$ to $\tau^{\prime}$, which is a strand in $\mathcal{V}$. Since level $L_{t}$ is not missing any nodes, then each of the nodes in $P_{\tau}$ has a unique child in the history tree, as well. It follows that all descendants of $\tau$ up to level $L_{t}$ have the same anonymity as $\tau$. Also, by definition of $t$ and the way it is updated (Line 14), no guesses have been made on nodes at levels deeper than $L_{t}$, and hence 1 is satisfied up to this point. Thus, Lemma A.4 applies.
@@
Observe that there are no counting cuts, or Line 13 would not be reachable. Due to Lines 9–11, all of the nodes in $P_{\tau}$ initially become guessers. Hence, all levels between $L_{s}$ and $L_{t-1}$ must have a non-empty set of guessers at all times. Consider any level $L_{i}$ with $s<i\leq t$ such that all the guessable nodes in $L_{i}$ are already counted. Let $S$ be the set of guessers in $L_{i-1}$; note that not all nodes in $L_{i-1}$ are guessers, or else they would give rise to a counting cut. Since the network is $1$-union-connected, there is a red edge $\{u,v\}$ (with positive multiplicity) such that $u\in S$ and the parent of $v$ is not in $S$. By definition, the node $v$ is guessable; therefore, it is counted. Also, since the parent of $v$ is not a guesser, $v$ must have a non-counted parent or a non-counted sibling; note that such a non-counted node is in $\mathcal{V}$.
@@
We have proved that every level between $L_{s+1}$ and $L_{t}$ lacking a guessable non-counted node contains a counted node $v$ having a parent or a sibling that is not counted: we call such a node $v$ a bad node. To conclude the proof, it suffices to show that there are at most $n-2$ bad nodes between $L_{s+1}$ and $L_{t}$. Observe that no nodes in $P_{\tau}$ can be bad.
@@
We will prove by induction that, if a subtree $\mathcal{W}$ of $\mathcal{V}$ contains the root $r$, the leader node $\tau$, no counting cuts, and no non-trivial isles, then $\mathcal{W}$ contains at most $f-1$ bad nodes, where $f$ is the number of leaves of $\mathcal{W}$ not in the subtree hanging from $\tau$. The case $f=0$ is impossible, because the single node $\tau$ yields a counting cut. Thus, the base case is $f=1$, which holds because any bad node $v$ in $\mathcal{W}$ and not in $P_{\tau}$ gives rise to the counting cut $\{\tau,v\}$ (recall that a bad node is counted by definition).
@@
For the induction step, let $v$ be a bad node of maximum depth in $\mathcal{W}$. Let $(v_{1},v_{2},\dots,v_{k})$ be the black path from $v_{1}=v$ to the root $v_{k}=r$, and let $1<i\leq k$ be the smallest index such that $v_{i}$ has more than one child in $\mathcal{W}$ ($i$ must exist, because this path eventually joins the black path from $\tau$ to $r$). Let $\mathcal{W}^{\prime}$ be the tree obtained by deleting the black edge $\{v_{i-1},v_{i}\}$ from $\mathcal{W}$, as well as the subtree hanging from it. Notice that the induction hypothesis applies to $\mathcal{W}^{\prime}$: since $v_{1}$ is counted, and each of the nodes $v_{2}$, …, $v_{i-1}$ has a unique child in $\mathcal{W}$, the removal of $\{v_{i-1},v_{i}\}$ does not create counting cuts or non-trivial isles. Also, $v_{2}$ is not counted (unless perhaps $v_{2}=v_{i}$), because $v_{1}$ is bad. Furthermore, none of the nodes $v_{3}$, …, $v_{i-1}$ is counted, or else $v_{2}$ would be an internal node of a (non-trivial) isle in $\mathcal{W}$. Therefore, $\mathcal{W}^{\prime}$ has exactly one less bad node than $\mathcal{W}$ and one less leaf; the induction hypothesis now implies that $\mathcal{W}$ contains at most $f-1$ bad nodes.
@@
Observe that the subtree $\mathcal{V}^{\prime}$ of $\mathcal{V}$ formed by all levels up to $L_{t}$ satisfies all of the above conditions, as it contains $\tau\in L_{s}$, the root $r$, and has no counting cuts, because a counting cut for $\mathcal{V}^{\prime}$ would be a counting cut for $\mathcal{V}$, as well (recall that $\mathcal{V}$ has no counting cuts). Also, Lemma A.4 (iii) ensures that $\mathcal{V}^{\prime}$ contains no non-trivial complete isles. However, since no nodes are missing from the levels of $\mathcal{V}^{\prime}$, all isles in $\mathcal{V}^{\prime}$ are complete, and thus must be trivial. We conclude that, if $\mathcal{V}^{\prime}$ has $f$ leaves not in the subtree hanging from $\tau$, it contains at most $f-1$ bad nodes. Since such leaves induce a partition of the at most $n-1$ processes not represented by $\tau$, we have $f\leq n-1$, implying that the number of bad nodes up to $L_{t}$ is at most $n-2$. ∎
Main lemma. We are now ready to prove Lemma 4.2.
Lemma A.7.
@@
Let $\texttt{ApproxCount}(\mathcal{V},s,x,\ell)$ return $(n^{\prime},t)$. Assume that $\tau$ exists and $x\geq a(\tau)$. Let $\tau^{\prime}$ be the (unique) descendant of $\tau$ in $\mathcal{V}$ at level $L_{t}$, and let $L_{t^{\prime}}$ be the last level of $\mathcal{V}$. Then:
@@
    If $x=a(\tau)=a(\tau^{\prime})$, then $n^{\prime}\neq-3$.
@@
    If $n^{\prime}>0$ and $t^{\prime}\geq t+n^{\prime}$ and $a(\tau)=a(\tau^{\prime})$, then $n^{\prime}=n$.
@@
    If $t^{\prime}\geq s+(\ell+2)n-1$, then $s\leq t\leq s+(\ell+1)n-1$ and $n^{\prime}\neq-1$. Moreover, if $n^{\prime}=-2$, then $L_{t}$ contains a leader node with at least two children in $\mathcal{V}$.
@@
Note that $\tau^{\prime}$ is well defined, because the returned pair is $(n^{\prime},t)$, which means that either $t=s$, and thus $\tau=\tau^{\prime}$, or $t>s$, and hence some guesses have been made on level $L_{t}$, the first of which must have had the parent of $\tau^{\prime}$ as the guesser.
@@
Let us prove statement (i). The assumption $x=a(\tau)$ implies $\delta=1$. Moreover, since $a(\tau)=a(\tau^{\prime})$, 1 is satisfied whenever a guess is made (this is a straightforward induction). Therefore, by Lemma A.4 (iv), all nodes marked as counted up to $L_{t}$ indeed have the correct guesses. So, the conditional anonymity that is computed for any node is equal to its anonymity ($a^{\prime}(v)=\delta a(v)=a(v)$), and hence is an integer. This implies that ApproxCount cannot return the error code $-3$ at Line 19. Also, either $\ell^{\prime}=\ell$ if all leader processes have been counted, or $\ell^{\prime}<\ell$ if some leader nodes are missing from the view. Either way, ApproxCount cannot return the error code $-3$ at Line 31. We conclude that $n^{\prime}\neq-3$.
@@
Let us prove statement (ii). Again, because $a(\tau)=a(\tau^{\prime})$, 1 is satisfied, and all nodes marked as counted have correct guesses. Also, $x\geq a(\tau)$ is equivalent to $\delta\geq 1$. By assumption, ApproxCount returns $(n^{\prime},t)$ with $n^{\prime}>0$ and $t^{\prime}\geq t+n^{\prime}$. Since $n^{\prime}>0$, a counting cut $C$ was found whose nodes are within levels up to $L_{t}$, and $n^{\prime}$ is the sum of the conditional anonymities of all nodes in $C$. Let $S_{C}$ be the set of processes represented by the nodes of $C$; note that $n^{\prime}\geq|S_{C}|$, because $\delta\geq 1$. We will prove that $S_{C}$ includes all processes in the system. Assume the contrary; [24, Lemma 4.2] implies that, since $t^{\prime}\geq t+n^{\prime}\geq t+|S_{C}|$, there is a node $z\in L_{t}$ representing some process not in $S_{C}$. Thus, the black path from $z$ to the root $r$ does not contain any node of $C$, contradicting the fact that $C$ is a counting cut with no nodes after $L_{t}$. Therefore, $|S_{C}|=n$, i.e., the nodes in $C$ represent all processes in the system. Since ApproxCount returns $n^{\prime}>0$, the “reality check” $\ell^{\prime}=\ell$ succeeds (Lines 30–32). However, $\ell^{\prime}$ is the sum of the conditional anonymities of all leader nodes in $C$, and hence $\ell^{\prime}=\delta\ell$, implying that $\delta=1$. Thus, $n^{\prime}=\delta n=n$, as claimed.
@@
Let us prove statement (iii). Once again, $x\geq a(\tau)$ is equivalent to $\delta\geq 1$. By [24, Corollary 4.3], if $L_{t^{\prime}}$ is the last level of $\mathcal{V}$, then no nodes are missing from level $L_{t^{\prime}-n}$. In particular, since $t^{\prime}-n\leq(\ell+1)n-1$, no nodes are missing from any level up to $L_{(\ell+1)n-1}$. Let $\tau^{\prime\prime}$ be the deepest descendant of $\tau$ that is marked as counted at Lines 9–11, and let $L_{p}$ be the level of $\tau^{\prime\prime}$. By construction, either all children of $\tau^{\prime\prime}$ are missing from $L_{p+1}$ or at least two children of $\tau^{\prime\prime}$ are in $L_{p+1}$. Also note that $\tau^{\prime}$ must be an ancestor of $\tau^{\prime\prime}$, and so $t\leq p$.
@@
Assume that $p<s+(\ell+1)n-1$. This implies that no nodes are missing from level $L_{p+1}$, and therefore $\tau^{\prime\prime}$ must have at least two children in $L_{p+1}$. Since $t\leq p$, we have $t<s+(\ell+1)n-1$, as desired. Now assume that $n^{\prime}=-2$, which implies that the algorithm was unable to find a counting cut. We claim that in this case $t=p$. So, assume for a contradiction that $t\leq p-1$. It follows that $\tau^{\prime}\in L_{t}$ is a guesser. Recall that $L_{t}$ and $L_{t+1}$ are not missing any nodes, because $t\leq p$. Since the network is connected at round $t+1$, there is at least one node in $L_{t+1}$ that is guessable by $\tau^{\prime}$, and so $L_{t+1}$ is a guessable level. However, the algorithm cannot return $n^{\prime}=-2$ as long as there are guessable levels (Line 12). Thus, $t=p$, which means that $\tau^{\prime}=\tau^{\prime\prime}$, and hence $\tau^{\prime}$ has at least two children in $\mathcal{V}$, as desired.
@@
Assume now that $p\geq s+(\ell+1)n-1$, and recall that no nodes are missing from the levels up to $L_{(\ell+1)n-1}$. In particular, since $\delta\leq\ell$, no nodes are missing from the levels in the interval $\mathcal{L}$, which consists of the $(\delta+1)n-1$ levels from $L_{s+1}$ to $L_{s+(\delta+1)n-1}$. Thus, by definition of $p$, as long as no guesses are made outside of $\mathcal{L}$, 1 holds, and therefore Lemmas A.5 and A.6 apply. Hence, as long as no guesses are made outside of $\mathcal{L}$, at most $\delta n$ levels of $\mathcal{L}$ are locked (Lemma A.5) and at most $n-2$ levels of $\mathcal{L}$ lack a guessable non-counted node (Lemma A.6). We conclude that $\mathcal{L}$ always contains at least one guessable level, and no guesses are ever made outside of $\mathcal{L}$ until either a counting cut is found or $n^{\prime}=-3$ is returned at Line 19. In both cases, $n^{\prime}=-2$ is not returned, and moreover $t\leq s+(\delta+1)n-1\leq s+(\ell+1)n-1$, as desired.

### 05 Apr 2024 v6

@@ 32   Else return $(n^{\prime},t)$
Note that the assumptions of Lemma 5.6 are satisfied, and therefore all the conditional anonymities and weights assigned to nodes up to this point are correct according to our definitions.
We will begin by proving that, if the subtree hanging from a node $v$ of $\mathcal{V}$ contains more than $a^{\prime}(v)$ nodes marked as guessed, then it contains a node $v^{\prime}$ marked as guessed such that $w(v^{\prime})>a^{\prime}(v^{\prime})$. The proof is by well-founded induction based on the subtree relation in $\mathcal{V}$. If $v$ is guessed, then we can take $v^{\prime}=v$, for in this case $w(v)>a^{\prime}(v)$. Otherwise, by the pigeonhole principle, $v$ has at least one child $u$ whose hanging subtree contains more than $a^{\prime}(u)$ guessed nodes. Thus, $v^{\prime}$ is found in this subtree by the induction hypothesis.
Now, assume for a contradiction that more than $\delta(n-1)$ levels of $\mathcal{V}$ contain locked nodes; in particular, $\mathcal{V}$ contains more than $\delta(n-1)$ nodes marked as guessed. Consider the nodes in level $L_{s}$ other than $\vartheta$; the sum of their anonymities is at most $n-a(\vartheta)$ (note that some nodes may be missing from $L_{s}$), and so the sum of their conditional anonymities is at most $\delta(n-a(\vartheta))\leq\delta(n-1)$. Thus, by the pigeonhole principle, there is a a node $v\neq\vartheta$ in $L_{s}$ whose hanging subtree contains more than $a^{\prime}(v)$ nodes marked as guessed.
Therefore, as proved above, the subtree hanging from $v$ contains a guessed node $v^{\prime}$ such that $w(v^{\prime})>a^{\prime}(v^{\prime})$. Since $\delta\geq 1$ and Lemma 5.6 (i) holds, we can apply Lemma 5.4 to $v^{\prime}$, which implies that there exist heavy nodes. In turn, this contradicts Lemma 5.6 (ii). We conclude that at most $\delta(n-1)$ levels contain locked nodes. ∎
Lemma 5.8.
Assume that $\delta\geq 1$. Then, as long as level $L_{t}$ of $\mathcal{V}$ is not missing any nodes (where $t$ is defined and updated as in ApproxCount), whenever Line 13 is reached, there are at most $n-2$ levels in the range from $L_{s+1}$ to $L_{t}$ where all guessable nodes are already counted.
@@
By definition of $t$, either $t=s$ or the algorithm has performed at least one guess on a node at level $L_{t}$ with a guesser at level $L_{t-1}$. It is easy to prove by induction that the first guesser to perform a guess on this level must be the unique descendant $\vartheta^{\prime}\in L_{t-1}$ of the selected leader node $\vartheta\in L_{s}$. Moreover, both $\vartheta^{\prime}$ and its unique child in $\mathcal{V}$ have been assigned conditional anonymity $x$ at Lines 9–11, and the same is true of all nodes in the black path $P_{\vartheta}$ from $\vartheta$ to $\vartheta^{\prime}$, which is a strand in $\mathcal{V}$. Since level $L_{t}$ is not missing any nodes, then each of the nodes in $P_{\vartheta}$ has a unique child in the history tree, as well. It follows that all descendants of $\vartheta$ up to level $L_{t}$ have the same anonymity as $\vartheta$. Also, by definition of $t$ and the way it is updated (Line 14), no guesses have been made on nodes at levels deeper than $L_{t}$, and hence 1 is satisfied up to this point. Thus, Lemma 5.6 applies.
@@
Observe that there are no counting cuts, or Line 13 would not be reachable. Due to Lines 9–11, all of the nodes in $P_{\vartheta}$ initially become guessers. Hence, all levels between $L_{s}$ and $L_{t-1}$ must have a non-empty set of guessers at all times. Consider any level $L_{i}$ with $s<i\leq t$ such that all the guessable nodes in $L_{i}$ are already counted. Let $S$ be the set of guessers in $L_{i-1}$; note that not all nodes in $L_{i-1}$ are guessers, or else they would give rise to a counting cut. Since the network is 1-union-connected, there is a red edge $\{u,v\}$ (with positive multiplicity) such that $u\in S$ and the parent of $v$ is not in $S$. By definition, the node $v$ is guessable; therefore, it is counted. Also, since the parent of $v$ is not a guesser, $v$ must have a non-counted parent or a non-counted sibling; note that such a non-counted node is in $\mathcal{V}$, because the levels up to $L_{t}$ are not missing any nodes.
@@
We have proved that every level between $L_{s+1}$ and $L_{t}$ where all guessable nodes are counted contains a counted node $v$ having a parent or a sibling that is not counted: we call such a node $v$ a bad node. To conclude the proof, it suffices to show that there are at most $n-2$ bad nodes between $L_{s+1}$ and $L_{t}$. Observe that no nodes in $P_{\vartheta}$ can be bad.
@@
We will prove that, if a subtree $\mathcal{W}$ of $\mathcal{V}$ contains the root $r$, the leader node $\vartheta$, no counting cuts, and no non-trivial isles, then $\mathcal{W}$ contains at most $f-1$ bad nodes, where $f$ is the number of leaves of $\mathcal{W}$ not in the subtree hanging from $\vartheta$. We stress that, in the context of $\mathcal{W}$, a bad node is defined as a counted node in $\mathcal{W}$ (other than $\vartheta$ and its descendants) that has a non-counted parent or a non-counted sibling in $\mathcal{W}$.
The proof is by induction on $f$. The case $f=0$ is impossible, because the single node $\vartheta$ yields a counting cut. Thus, the base case is $f=1$, which holds because any bad node $v$ in $\mathcal{W}$ and not in $P_{\vartheta}$ gives rise to the counting cut $\{\vartheta,v\}$ (recall that a bad node is counted by definition).
@@
For the induction step, let $v$ be a bad node of maximum depth in $\mathcal{W}$. Let $(v_{1},v_{2},\dots,v_{k})$ be the black path from $v_{1}=v$ to the root $v_{k}=r$, and let $1<i\leq k$ be the smallest index such that $v_{i}$ has more than one child in $\mathcal{W}$ ($i$ must exist, because this path eventually joins the black path from $\vartheta$ to $r$). Let $\mathcal{W}^{\prime}$ be the tree obtained by deleting the black edge $\{v_{i-1},v_{i}\}$ from $\mathcal{W}$, as well as the subtree hanging from it.
Notice that the induction hypothesis applies to $\mathcal{W}^{\prime}$: since $v_{1}$ is counted, and each of the nodes $v_{2}$, …, $v_{i-1}$ has a unique child in $\mathcal{W}$, the removal of $\{v_{i-1},v_{i}\}$ does not create counting cuts or non-trivial isles. Also, $v_{2}$ is not counted (unless $v_{2}=v_{i}$), because $v_{1}$ is bad. Furthermore, none of the nodes $v_{3}$, …, $v_{i-1}$ is counted, or else $v_{2}$ would be an internal node of a (non-trivial) isle in $\mathcal{W}$. Therefore, none of the nodes $v_{2}$, …, $v_{i-1}$ is counted. In particular, none of these nodes is bad in $\mathcal{W}$.
Moreover, a node of $\mathcal{W}^{\prime}$ is bad if and only if it is bad in $\mathcal{W}$. This is trivial for all nodes, except for the siblings of $v_{i-1}$, which require a careful proof. Let $u\neq v_{i-1}$ be a sibling of $v_{i-1}$ in $\mathcal{W}$. If $u$ is not counted, then it is not a bad node in $\mathcal{W}$ nor in $\mathcal{W}^{\prime}$. Hence, let us assume that $u$ is counted. If $v_{i}$ is not counted, then $u$ is a bad node in $\mathcal{W}$ and in $\mathcal{W}^{\prime}$. Thus, let us assume that $v_{i}$ is counted. Note that $i=2$ must hold, otherwise $v$ and $v_{i}$ would form a non-trivial isle in $\mathcal{W}$. In particular, $v$ and $u$ are siblings. Since $v$ is a bad node and its parent $v_{2}$ is counted, it must have a non-counted sibling $u^{\prime}\neq u$ in $\mathcal{W}$. However, $u^{\prime}$ is also a node of $\mathcal{W}^{\prime}$, and therefore $u$ is not a bad node in $\mathcal{W}$ nor in $\mathcal{W}^{\prime}$.
It follows that $\mathcal{W}^{\prime}$ has exactly one less bad node than $\mathcal{W}$ and at most $f-1$ leaves (because there is at least one leaf in the subtree of $\mathcal{W}$ hanging from $v$). Thus, the induction hypothesis implies that $\mathcal{W}^{\prime}$ contains at most $f-2$ bad nodes, and therefore $\mathcal{W}$ contains at most $f-1$ bad nodes.
@@
Observe that the subtree $\mathcal{V}^{\prime}$ of $\mathcal{V}$ formed by all levels up to $L_{t}$ satisfies all of the above conditions, as it contains $\vartheta\in L_{s}$, the root $r$, and has no counting cuts, because a counting cut for $\mathcal{V}^{\prime}$ would be a counting cut for $\mathcal{V}$, as well (recall that $\mathcal{V}$ has no counting cuts). Also, Lemma 5.6 (iii) ensures that $\mathcal{V}^{\prime}$ contains no non-trivial complete isles. However, since no nodes are missing from the levels of $\mathcal{V}^{\prime}$, all isles in $\mathcal{V}^{\prime}$ are complete, and thus must be trivial. We conclude that, if $\mathcal{V}^{\prime}$ has $f$ leaves not in the subtree hanging from $\vartheta$, it contains at most $f-1$ bad nodes. Since such leaves induce a partition of the at most $n-1$ processes not represented by $\vartheta$, we have $f\leq n-1$, implying that the number of bad nodes up to $L_{t}$ is at most $n-2$. ∎
Observe that the statement of Lemma 5.8 holds for $n=1$ as well, because in this case the single node $\vartheta$ constitutes a counting cut, and Line 13 is never reached.
Main lemma. We are now ready to prove the salient properties of ApproxCount as summarized in Lemma 5.1, which we restate next.
Lemma 5.9.
@@
Let $\texttt{ApproxCount}(\mathcal{V},s,x,\ell)$ return $(n^{\prime},t)$. Assume that $\vartheta$ exists and $x\geq a(\vartheta)$. Let ϑsuperscriptitalic-$\vartheta^{\prime}$ be the (unique) descendant of $\vartheta$ in $\mathcal{V}$ at level $L_{t}$, and let $L_{t^{\prime}}$ be the last level of $\mathcal{V}$. Then:
@@
    If $x=a(\vartheta)=a(\vartheta^{\prime})$, then $n^{\prime}\neq-3$.
@@
    If $n^{\prime}>0$ and $t^{\prime}\geq t+n^{\prime}$ and $a(\vartheta)=a(\vartheta^{\prime})$, then $n^{\prime}=n$.
@@
    If $t^{\prime}\geq s+(\ell+2)(n-1)$, then $s\leq t\leq s+(\ell+1)(n-1)$ and $n^{\prime}\neq-1$. Moreover, if $n^{\prime}=-2$, then $L_{t}$ contains a leader node with at least two children in $\mathcal{V}$.
@@
Note that $\vartheta^{\prime}$ is well defined, because the returned pair is $(n^{\prime},t)$, which means that either $t=s$, and thus $\vartheta=\vartheta^{\prime}$, or $t>s$, and hence some guesses have been made on nodes in level $L_{t}$, the first of which must have had the parent of $\vartheta^{\prime}$ as the guesser.
@@
Let us prove statement (i). The assumption $x=a(\vartheta)$ implies $\delta=1$. Moreover, since $a(\vartheta)=a(\vartheta^{\prime})$, 1 is satisfied whenever a guess is made (this is a straightforward induction). Therefore, by Lemma 5.6 (iv), all nodes marked as counted up to $L_{t}$ indeed have the correct guesses. So, the conditional anonymity that is computed for any node is equal to its anonymity ($a^{\prime}(v)=\delta a(v)=a(v)$), and hence is an integer. This implies that ApproxCount cannot return the error code $-3$ at Line 19. Also, either $\ell^{\prime}=\ell$ if all leader processes have been counted, or $\ell^{\prime}<\ell$ if some leader nodes are missing from the view. Either way, ApproxCount cannot return the error code $-3$ at Line 31. We conclude that $n^{\prime}\neq-3$.
@@
Let us prove statement (ii). Again, because $a(\vartheta)=a(\vartheta^{\prime})$, 1 is satisfied, and all nodes marked as counted have correct guesses. Also, $x\geq a(\vartheta)$ is equivalent to $\delta\geq 1$. By assumption, ApproxCount returns $(n^{\prime},t)$ with $n^{\prime}>0$ and $t^{\prime}\geq t+n^{\prime}$. Since $n^{\prime}>0$, a counting cut $C$ was found whose nodes are within levels up to $L_{t}$, and $n^{\prime}$ is the sum of the conditional anonymities of all nodes in $C$. Let $S_{C}$ be the set of processes represented by the nodes of $C$; note that $n^{\prime}\geq|S_{C}|$, because $\delta\geq 1$. We will prove that $S_{C}$ includes all processes in the system. Assume the contrary; Lemma 3.3 implies that, since $t^{\prime}\geq t+n^{\prime}\geq t+|S_{C}|$, there is a node $z\in L_{t}$ representing some process not in $S_{C}$. Thus, the black path from $z$ to the root $r$ does not contain any node of $C$, contradicting the fact that $C$ is a counting cut with no nodes after $L_{t}$. Therefore, $|S_{C}|=n$, i.e., the nodes in $C$ represent all processes in the system. Since ApproxCount returns $n^{\prime}>0$, the “reality check” $\ell^{\prime}=\ell$ succeeds (Lines 30–32). However, $\ell^{\prime}$ is the sum of the conditional anonymities of all leader nodes in $C$, and hence $\ell^{\prime}=\delta\ell$, implying that $\delta=1$. Thus, $n^{\prime}=\delta n=n$, as claimed.
@@
Let us prove statement (iii). Once again, $x\geq a(\vartheta)$ is equivalent to $\delta\geq 1$. By Corollary 3.4, if $L_{t^{\prime}}$ is the last level of $\mathcal{V}$, then no nodes are missing from level $L_{t^{\prime}-n+1}$. In fact, since $t^{\prime}-n+1\geq s+(\ell+1)(n-1)$, no nodes are missing from any level up to $L_{s+(\ell+1)(n-1)}$. Let $\vartheta^{\prime\prime}$ be the deepest descendant of $\vartheta$ that is marked as counted at Lines 9–11, and let $L_{p}$ be the level of $\vartheta^{\prime\prime}$. By construction, either all children of $\vartheta^{\prime\prime}$ are missing from $L_{p+1}$ or at least two children of $\vartheta^{\prime\prime}$ are in $L_{p+1}$. Also note that $\vartheta^{\prime}$ must be an ancestor of $\vartheta^{\prime\prime}$, and so $t\leq p$.
@@
Assume that $p<s+(\ell+1)(n-1)$. This implies that no nodes are missing from level $L_{p+1}$, and therefore $\vartheta^{\prime\prime}$ must have at least two children in $L_{p+1}$. Since $t\leq p$, we have $t<s+(\ell+1)(n-1)$, as desired. Now assume that $n^{\prime}=-2$, which implies that the algorithm was unable to find a counting cut. We claim that in this case $t=p$. So, assume for a contradiction that $t\leq p-1$. It follows that $\vartheta^{\prime}\in L_{t}$ is a guesser. Recall that $L_{t}$ and $L_{t+1}$ are not missing any nodes, because $t\leq p$. Since the network is connected at round $t+1$, there is at least one node $v\in L_{t+1}$ that is guessable by $\vartheta^{\prime}$. Also, since no guess has ever been made in level $L_{t+1}$ (due to the way $t$ is updated in Line 14), it follows that $v$ is not counted and not locked. However, the algorithm cannot return $n^{\prime}=-2$ as long as there are nodes such as $v$ (Line 12). Thus, $t=p$, which means that $\vartheta^{\prime}=\vartheta^{\prime\prime}$, and hence $\vartheta^{\prime}$ has at least two children in $\mathcal{V}$, as desired.
@@
Assume now that $p\geq s+(\ell+1)(n-1)$. If $\vartheta$ is the only node in $L_{s}$, it constitutes a counting cut. In this special case, Line 13 is never reached, $n^{\prime}=-2$ is not returned at Line 25, and $t=s$. Hence, we may assume that there are nodes in $L_{s}$ other than $\vartheta$ and Line 13 is reached (so, we have $n>1$). Recall that no nodes are missing from any level up to $L_{s+(\ell+1)(n-1)}$. In particular, no nodes are missing from the levels in the non-empty interval $\mathcal{L}$ consisting of the $(\ell+1)(n-1)$ levels from $L_{s+1}$ to $L_{s+(\ell+1)(n-1)}$. Thus, by definition of $p$, as long as no guesses are made outside of $\mathcal{L}$, 1 holds, and therefore Lemmas 5.7 and 5.8 apply. Hence, as long as no guesses are made outside of $\mathcal{L}$, there are at most $\delta(n-1)$ levels of $\mathcal{L}$ containing locked nodes (Lemma 5.7) and there are at most $n-2$ levels of $\mathcal{L}$ where all guessable nodes are already counted (Lemma 5.8). So, there are at most
$\delta(n-1)+n-2\leq\ell(n-1)+n-2=(\ell+1)(n-1)-1$
levels in $\mathcal{L}$ were the algorithm can make no new guesses. We conclude that $\mathcal{L}$ always contains at least one node where a guess can be made, and no guesses are ever made outside of $\mathcal{L}$ until either a counting cut is found or $n^{\prime}=-3$ is returned at Line 19. In both cases, $n^{\prime}=-2$ is not returned, and moreover $t\leq s+(\ell+1)(n-1)$, as desired.